OpenStudy (anonymous):

Given the general identity tan X =sin X/cos X , which equation relating the acute angles, A and C, of a right ∆ABC is true?

OpenStudy (anonymous):

A. tan A = sin A/sin C B. cos A = tan (90-A)/sin (90-C) C. sin C = cos A/tan C D. cos A = tan C E. sin C = cos (90-C)/tan A

OpenStudy (amorfide):

okay so let's draw a triangle

OpenStudy (amorfide):